# Unified Multimodal Understanding and Generation Models: Advances, Challenges, and Opportunities

### v4

@@ 2.2 Text-to-Image Model
Nexus-Gen [163] b-3 Qwen2.5-VL QwenViT Learnable Query FLUX-1.dev Causal 2025-04
@@
OpenUni [166] b-3 InternVL3 SigLIP Learnable Query SANA-1.5 Causal 2025-05
@@
[…]
@@ 3.2 Auto-Regressive Models
2) Semantic Encoding. To overcome the semantic limitations inherent in pixel-based encoders, a growing body of work adopts semantic encoding, where image inputs are processed using pretrained text-aligned vision encoders such as OpenAI-CLIP [22], SigLIP [202], EVA-CLIP [36], or more recent unified tokenizers like UNIT [216], as shown in Fig. 5 (b-2). Some of these models leverage the multimodal features encoded by the multimodal autoregressive model as conditions for a diffusion model, enabling image generation with retained multimodal understanding capabilities, like OmniGen2 [158] that utilizes Qwen2.5-VL [10] as the multimodal model and enhanced OmniGen [217] as the image diffusion model. However, most of these models are trained on large-scale image-text pairs with contrastive or regression-based objectives, producing visual embeddings that align closely with language features in a shared semantic space. Such representations enable more effective cross-modal alignment and are particularly beneficial for multimodal understanding and generation.
@@
Several representative models leverage different semantic encoders and architectural designs to support unified multimodal tasks. Emu [142], Emu2 [33], and LaViT [143] all employ EVA-CLIP [36] as their vision encoder. Notably, Emu [142] introduces the initial architecture combining a frozen EVA-CLIP encoder, a large language model, and a diffusion decoder to unify VQA, image captioning, and image generation. Emu2 [33] builds upon Emu [142] by proposing a simplified and scalable modeling framework for unified multimodal pretraining. It scales the MLLM model up to 37B parameters, significantly enhancing both understanding and generation capabilities. LaViT [143] introduces a dynamic visual tokenization mechanism built on top of EVA-CLIP. It employs a selector and merger module to adaptively select visual tokens from image embeddings based on content complexity. This process dynamically determines the length of the visual token sequence per image. The dynamic tokenization significantly reduces redundant information while preserving important visual cues, improving training efficiency and generation quality in tasks such as captioning, visual question answering, and image generation. DreamLLM [34], VL-GPT [35], and MM-Interleaved [144], and PUMA [147] utilize OpenAI-CLIP encoder [22]. DreamLLM [34] introduces a lightweight linear projection to align CLIP embeddings with language tokens, while VL-GPT [35] employs a powerful casual transformer after OpenAI-CLIP vision encoder to effectively retain both semantic information and pixel details of the original image. Both MM-Interleaved [144] and PUMA [147] extract multi-granular image features via a CLIP tokenizer with simple ViT-Adapter or pooling operation to provide fine-grained feature fusion, thus supporting rich multimodal generation. Mini-Gemini [145] introduces a visual token enhancement mechanism that requires dual semantic encoders. Specifically, it leverages a CLIP-pretrained ViT encoder [22] to obtain global visual tokens, while a LAION-pretrained ConvNeXt encoder provides dense local visual information. A cross-attention module is then employed to refine the global visual tokens by incorporating detailed visual cues from the dense encoder. These enhanced global tokens are subsequently combined with text tokens and processed by an LLM for joint vision-language understanding and generation. This design effectively bridges the semantic abstraction of CLIP features with the pixel-level precision of dense encoders. MetaMorph [148] employs SigLIP [202] to extract visual embeddings and introduces modality-specific adapters within a pretrained language model. These adapters are inserted throughout multiple transformer layers, allowing for deeper vision-language interaction compared to shallow projection approaches. ILLUME [149] adopt UNIT [216] as its vision encoder to provide a unified representation that balances semantic alignment and pixel-level fidelity. Unlike CLIP-like encoders that focus purely on contrastive objectives, UNIT [216] is jointly trained with both image reconstruction and contrastive alignment losses, producing tokens suitable for both vision-language understanding and image synthesis. Built on the powerful UNIT tokenizer, ILLUME effectively generates image tokens that retain both semantic and pixel-level information, which achieves better performance in multiple understanding and generation tasks, including captioning, VQA, text-to-image, and interleaved generation. Similarly, VILA-U [146] and UniTok [150] mimic UNIT [216] to introduce image-text contrastive learning to obtain a novel text-aligned vision tokenizer that balances semantic alignment and pixel-level fidelity. QLIP [151] addresses the potential conflict between reconstruction and text-image alignment tasks by implementing binary-spherical quantization. Tar [157] initiates the visual codebook by leveraging the vocabulary of LLMs and incorporates scale-adaptive pooling and decoding methodologies. This approach enables the model to adjust the length of the tokenizer according to the requirement: employing coarse-grained tokenizers for efficient generation and fine-grained tokenizers for comprehensive understanding. In generation tasks, Tar utilizes diffusion techniques to enhance the visual generation outcomes of AR models. UniFork [153] capitalizes on the text-aligned vision features of VILA-U. However, differentiating itself from the fully-shared parameters of understanding and generation MLLM, UniFork shares the parameters solely with these tasks at the shallow layer. At the deeper layer, these tasks are managed by distinct networks. This architecture successfully mediates the equilibrium between shared learning and task-specific specialization. UniCode2 [154] employs a cascaded codebook. In line with the method outlined in [218], it utilizes a substantial codebook derived from clustered SigLIP feature as the frozen foundational codebook, while introducing supplementary learnable codebooks to refine semantics specific to particular tasks. This separation enhances utilization and fosters robust learning. Recent work DualToken [152] uses shallow-layer features of SigLIP for reconstruction and deep-layer features of SigLIP for semantic learning, thereby obtaining the texture and semantic visual features simultaneously. As a result, DualToken [152] achieves superior performance in both reconstruction and semantic tasks while demonstrating remarkable effectiveness in downstream MLLM understanding and generation tasks.
@@
Current implementations of learnable query encoding can be broadly divided into two representative paradigms. The first is represented by SEED [159], which proposes a seed tokenizer that learns causal visual embeddings. Specifically, an input image is first encoded into dense token features via a BLIP-2 ViT encoder [53]. These features are then concatenated with a set of learnable query tokens and processed by a causal Q-Former to produce causal visual embeddings. This design is trained using both image-text contrastive learning and image reconstruction supervision, allowing the learned embeddings to simultaneously retain low-level visual detail and capture high-level semantic alignment with text. Building on this foundation, SEED-LLAMA [160] and SEED-X [161] enhance the model’s capacity by replacing the OPT backbone [222] with a stronger LLaMA2 model [2] and upgrading the decoder to UnCLIP-SD [14] or SDXL [219], leading to improved performance in both understanding and generation tasks. The second approach, introduced by MetaQueries [162], provides a simplified version of learnable query encoding. Here, image features are extracted via a frozen SigLIP encoder [202], which are then concatenated with learnable query tokens and directly passed through a frozen vision-language backbone such as LLaVA [209] or Qwen2.5-VL [10]. The output causal embeddings are used as conditioning inputs for a diffusion-based image decoder, enabling high-quality image generation. Because the backbone is kept frozen, the vision-language understanding capabilities remain consistent with the underlying pretrained models, offering a lightweight yet effective solution for multimodal generation. Open-Uni [166] refines the architecture of MetaQueries by utilizing solely learnable queries and a lightweight connector between a MLLM and a diffusion model, facilitating cohesive multimodal understanding and generation. Open-Uni demonstrates that the connetor between the MLLM visual understanding component and the diffusion-based visual generation component can be minimal in complexity, exemplified by a configuration comprising merely six Transformer layers. Nexus-Gen [163] and Ming-Lite-Uni [164] follow the MetaQueries paradigm, but with notable advancements to further enhance multimodal generation. Nexus-Gen [163] introduces a more powerful diffusion decoder, FLUX-1.dev, which significantly improves the generation quality. This approach allows the model to better capture the intricate details and high-fidelity features necessary for complex image generation tasks. On the other hand, Ming-Lite-Uni [164] takes a different route by introducing a highly capable MLLM model, M2-omini [193], for enhanced vision-language interaction. This model performs advanced vision-language conditioning to generate the conditioned image embeddings, ensuring a more semantically aligned representation. In addition, Ming-Lite-Uni fine-tunes its diffusion model by incorporating multi-scale learnable tokens, which facilitate improved semantic alignment across various visual scales. The multi-scale representation alignment mechanism enhances the model’s ability to generate detailed and contextually rich images from textual prompts, addressing challenges such as resolution mismatches and semantic inconsistencies. This innovative approach makes Ming-Lite-Uni a powerful tool for multimodal understanding and generation, pushing the boundaries of current methods in both flexibility and performance. Ming-Omni [167] adheres to the integrated MoE architecture, wherein modality-specific routing is facilitated through dedicated mechanisms tailored for each token, thereby enabling customized routing distributions. To address the multi-scale phenomenon inherent in visual generation [113], Ming-Omni employs multi-scale learnable queries, directed by an alignment strategy, to iteratively generate images progressing from coarse to fine detail. Furthermore, Ming-Omni integrates the audio modality and implements a dual-stage training strategy to mitigate the mutual influence between audio comprehension and generation tasks. The initial stage emphasizes comprehension capabilities, while the subsequent stage concentrates on enhancing generation quality. BLIP3-o [165] also employs learnable queries to bridge multimodal understanding and generation. However, it utilizes two diffusion models: one for learning CLIP embeddings and the other for using CLIP as a condition to generate images. It reveals that flow matching loss is more effective than MSE loss, enabling more diverse image sampling and yielding better image quality. To sum up, these learnable query-based designs share a common strength: they provide adaptive, compact, and semantically enriched representations that support both efficient image understanding and high-quality generation. By focusing on task-driven token extraction, such models offer a flexible and extensible alternative to traditional visual tokenizers, especially in unified multimodal frameworks.
@@
Pseudo Hybrid Encoding. Representative works in this category include Janus [168], Janus-Pro [169], OmniMamba [170], Unifluid [171], and MindOmni [172]. As shown in Fig. 5 (b-4), these models adopt dual encoders—typically a semantic encoder (e.g., SigLIP) and a pixel encoder (e.g., VQGAN or VAE)—but use them in a task-specific manner. During training, the semantic encoder branch is enabled for vision-language understanding tasks, while the pixel encoder branch is activated for image generation tasks. Although the dual encoders are trained concurrently with combined understanding and generation datasets, the pixel encoder is not utilized during inference in understanding tasks and the semantic encoder is disabled for text-to-image generation. However, for image editing, Unifluid [171] uses the semantic encoder to encode the source image while MindOmni [172] utilizes both VAE and semantic encoder to encode the source image. The rationale behind this design choice is that mixed training with both types of data can enhance performance across understanding and generation tasks. However, since only one encoder is active at any given time, these models do not fully harness the advantages of hybrid encoding. Specifically, they miss the opportunity to employ semantic grounding in generation tasks and fail to utilize high-fidelity visual details in comprehension tasks. Consequently, these models typically engage pixel decoders to reconstruct images from latent codes.
@@ 4.2 Text-to-Image Datasets
  • Synthesized Datasets: Specialized datasets for text-to-image synthesis are increasingly created using existing generative models. The text-to-image-2M dataset [254] provides  2 million enhanced text-image pairs for fine-tuning, curated using advanced T2I and captioning models. SFHQ-T2I [256] offers 122K diverse, high-resolution synthetic face images generated by multiple T2I models, ensuring variance and privacy. For entity control, the EliGen TrainSet [257] uses images from a baseline model (FLUX.1-dev) and MLLM-generated prompts for stylistic consistency and detailed annotation. Similarly, BLIP-3o 60k [165] provides 60,000 instruction tuning samples distilled from GPT-4o, covering various categories for diverse training. Lastly, ShareGPT-4o-Image[259] contributes 45K text-to-image pairs, where prompts are generated through both a structured attribute-first approach and an image-first approach, with corresponding images synthesized by GPT-4o’s image generation capabilities to distill its advanced skills.

### v5

@@ 2.2 Text-to-Image Model
Ovis-U1 [159] b-2 Ovis AimV2 MMDiT Causal 2025-06
X-Omni [160] b-2 Qwen2.5-VL QwenViT Siglip FLUX Causal 2025-07
Qwen-Image [161] b-2 Qwen2.5-VL QwenViT MMDiT Causal 2025-08
Bifrost-1 [162] b-2 Qwen2.5-VL QwenViT ViT FLUX Causal 2025-08
@@
Nexus-Gen [167] b-3 Qwen2.5-VL QwenViT Learnable Query FLUX Causal 2025-04
@@
OpenUni [170] b-3 InternVL3 InternViT Learnable Query Sana Causal 2025-05
@@
UniLIP [172] b-3 InternVL3 InternViT Learnable Query Sana Causal 2025-07
TBAC-UniImage [173] b-3 Qwen2.5-VL QwenViT Learnable Query Sana Causal 2025-08
@@
Skywork UniPic [179] b-4 Qwen2.5 SigLIP2 SDXL-VAE SDXL-VAE Causal 2025-08
MUSE-VL [180] b-5 Qwen-2.5&Yi-1.5 SigLIP VQGAN VQGAN Causal 2024-11
@@ 3.2 Auto-Regressive Models
2) Semantic Encoding. To overcome the semantic limitations inherent in pixel-based encoders, a growing body of work adopts semantic encoding, where image inputs are processed using pretrained text-aligned vision encoders such as OpenAI-CLIP [22], SigLIP [209], EVA-CLIP [36], or more recent unified tokenizers like UNIT [223], as shown in Fig. 5 (b-2). Some of these models leverage the multimodal features encoded by the multimodal autoregressive model as conditions for a diffusion model, enabling image generation with retained multimodal understanding capabilities, like OmniGen2 [158] that utilizes Qwen2.5-VL [10] as the multimodal model and enhanced OmniGen [224] as the image diffusion model, Ovis-U1 [159] extending the multimodal model Ovis [12] into a unified model by incorporating a custom-designed diffusion transformer while Qwen-Image [161] similarly building upon Qwen2.5-VL [10] by integrating a diffusion transformer. However, most of these models are trained on large-scale image-text pairs with contrastive or regression-based objectives, producing visual embeddings that align closely with language features in a shared semantic space. Such representations enable more effective cross-modal alignment and are particularly beneficial for multimodal understanding and generation.
@@
Several representative models leverage different semantic encoders and architectural designs to support unified multimodal tasks. Emu [142], Emu2 [33], and LaViT [143] all employ EVA-CLIP [36] as their vision encoder. Notably, Emu [142] introduces the initial architecture combining a frozen EVA-CLIP encoder, a large language model, and a diffusion decoder to unify VQA, image captioning, and image generation. Emu2 [33] builds upon Emu [142] by proposing a simplified and scalable modeling framework for unified multimodal pretraining. It scales the MLLM model up to 37B parameters, significantly enhancing both understanding and generation capabilities. Bifrost-1 [162] employs two semantic encoders, ViT for generation and the one employed in the used MLLM (QWen2.5-VL) for understanding. The predicted CLIP latents are used to bridge the MLLM and diffusion model. LaViT [143] introduces a dynamic visual tokenization mechanism built on top of EVA-CLIP. It employs a selector and merger module to adaptively select visual tokens from image embeddings based on content complexity. This process dynamically determines the length of the visual token sequence per image. The dynamic tokenization significantly reduces redundant information while preserving important visual cues, improving training efficiency and generation quality in tasks such as captioning, visual question answering, and image generation. DreamLLM [34], VL-GPT [35], and MM-Interleaved [144], and PUMA [147] utilize OpenAI-CLIP encoder [22]. DreamLLM [34] introduces a lightweight linear projection to align CLIP embeddings with language tokens, while VL-GPT [35] employs a powerful casual transformer after OpenAI-CLIP vision encoder to effectively retain both semantic information and pixel details of the original image. Both MM-Interleaved [144] and PUMA [147] extract multi-granular image features via a CLIP tokenizer with simple ViT-Adapter or pooling operation to provide fine-grained feature fusion, thus supporting rich multimodal generation. Mini-Gemini [145] introduces a visual token enhancement mechanism that requires dual semantic encoders. Specifically, it leverages a CLIP-pretrained ViT encoder [22] to obtain global visual tokens, while a LAION-pretrained ConvNeXt encoder provides dense local visual information. A cross-attention module is then employed to refine the global visual tokens by incorporating detailed visual cues from the dense encoder. These enhanced global tokens are subsequently combined with text tokens and processed by an LLM for joint vision-language understanding and generation. This design effectively bridges the semantic abstraction of CLIP features with the pixel-level precision of dense encoders. MetaMorph [148] employs SigLIP [209] to extract visual embeddings and introduces modality-specific adapters within a pretrained language model. These adapters are inserted throughout multiple transformer layers, allowing for deeper vision-language interaction compared to shallow projection approaches. ILLUME [149] adopt UNIT [223] as its vision encoder to provide a unified representation that balances semantic alignment and pixel-level fidelity. Unlike CLIP-like encoders that focus purely on contrastive objectives, UNIT [223] is jointly trained with both image reconstruction and contrastive alignment losses, producing tokens suitable for both vision-language understanding and image synthesis. Built on the powerful UNIT tokenizer, ILLUME effectively generates image tokens that retain both semantic and pixel-level information, which achieves better performance in multiple understanding and generation tasks, including captioning, VQA, text-to-image, and interleaved generation. Similarly, VILA-U [146] and UniTok [150] mimic UNIT [223] to introduce image-text contrastive learning to obtain a novel text-aligned vision tokenizer that balances semantic alignment and pixel-level fidelity. QLIP [151] addresses the potential conflict between reconstruction and text-image alignment tasks by implementing binary-spherical quantization. Tar [157] initiates the visual codebook by leveraging the vocabulary of LLMs and incorporates scale-adaptive pooling and decoding methodologies. This approach enables the model to adjust the length of the tokenizer according to the requirement: employing coarse-grained tokenizers for efficient generation and fine-grained tokenizers for comprehensive understanding. In generation tasks, Tar utilizes diffusion techniques to enhance the visual generation outcomes of AR models. UniFork [153] capitalizes on the text-aligned vision features of VILA-U. However, differentiating itself from the fully-shared parameters of understanding and generation MLLM, UniFork shares the parameters solely with these tasks at the shallow layer. At the deeper layer, these tasks are managed by distinct networks. This architecture successfully mediates the equilibrium between shared learning and task-specific specialization. UniCode2 [154] employs a cascaded codebook. In line with the method outlined in [225], it utilizes a substantial codebook derived from clustered SigLIP feature as the frozen foundational codebook, while introducing supplementary learnable codebooks to refine semantics specific to particular tasks. This separation enhances utilization and fosters robust learning. Recent work DualToken [152] uses shallow-layer features of SigLIP for reconstruction and deep-layer features of SigLIP for semantic learning, thereby obtaining the texture and semantic visual features simultaneously. As a result, DualToken [152] achieves superior performance in both reconstruction and semantic tasks while demonstrating remarkable effectiveness in downstream MLLM understanding and generation tasks. X-Omni [160] utilizes SigLIP-VQ as a visual encoder and employs reinforcement learning to mitigate the cumulative error associated with autoregressive inference and to reduce the information loss inherent in discrete encoding. This methodology substantially enhances the generation quality of discrete autoregressive models, facilitating a seamless integration of image and language generation.
@@
Current implementations of learnable query encoding can be broadly divided into two representative paradigms. The first is represented by SEED [163], which proposes a seed tokenizer that learns causal visual embeddings. Specifically, an input image is first encoded into dense token features via a BLIP-2 ViT encoder [53]. These features are then concatenated with a set of learnable query tokens and processed by a causal Q-Former to produce causal visual embeddings. This design is trained using both image-text contrastive learning and image reconstruction supervision, allowing the learned embeddings to simultaneously retain low-level visual detail and capture high-level semantic alignment with text. Building on this foundation, SEED-LLAMA [164] and SEED-X [165] enhance the model’s capacity by replacing the OPT backbone [229] with a stronger LLaMA2 model [2] and upgrading the decoder to UnCLIP-SD [14] or SDXL [226], leading to improved performance in both understanding and generation tasks. The second approach, introduced by MetaQueries [166], provides a simplified version of learnable query encoding. Here, image features are extracted via a frozen SigLIP encoder [209], which are then concatenated with learnable query tokens and directly passed through a frozen vision-language backbone such as LLaVA [216] or Qwen2.5-VL [10]. The output causal embeddings are used as conditioning inputs for a diffusion-based image decoder, enabling high-quality image generation. Because the backbone is kept frozen, the vision-language understanding capabilities remain consistent with the underlying pretrained models, offering a lightweight yet effective solution for multimodal generation. Open-Uni [170] refines the architecture of MetaQueries by utilizing solely learnable queries and a lightweight connector between a MLLM and a diffusion model, facilitating cohesive multimodal understanding and generation. Open-Uni demonstrates that the connetor between the MLLM visual understanding component and the diffusion-based visual generation component can be minimal in complexity, exemplified by a configuration comprising merely six Transformer layers. Nexus-Gen [167] and Ming-Lite-Uni [168] follow the MetaQueries paradigm, but with notable advancements to further enhance multimodal generation. Nexus-Gen [167] introduces a more powerful diffusion decoder, FLUX-1.dev, which significantly improves the generation quality. This approach allows the model to better capture the intricate details and high-fidelity features necessary for complex image generation tasks. On the other hand, Ming-Lite-Uni [168] takes a different route by introducing a highly capable MLLM model, M2-omini [200], for enhanced vision-language interaction. This model performs advanced vision-language conditioning to generate the conditioned image embeddings, ensuring a more semantically aligned representation. In addition, Ming-Lite-Uni fine-tunes its diffusion model by incorporating multi-scale learnable tokens, which facilitate improved semantic alignment across various visual scales. The multi-scale representation alignment mechanism enhances the model’s ability to generate detailed and contextually rich images from textual prompts, addressing challenges such as resolution mismatches and semantic inconsistencies. This innovative approach makes Ming-Lite-Uni a powerful tool for multimodal understanding and generation, pushing the boundaries of current methods in both flexibility and performance. Ming-Omni [171] adheres to the integrated MoE architecture, wherein modality-specific routing is facilitated through dedicated mechanisms tailored for each token, thereby enabling customized routing distributions. To address the multi-scale phenomenon inherent in visual generation [113], Ming-Omni employs multi-scale learnable queries, directed by an alignment strategy, to iteratively generate images progressing from coarse to fine detail. Furthermore, Ming-Omni integrates the audio modality and implements a dual-stage training strategy to mitigate the mutual influence between audio comprehension and generation tasks. The initial stage emphasizes comprehension capabilities, while the subsequent stage concentrates on enhancing generation quality. BLIP3-o [169] also employs learnable queries to bridge multimodal understanding and generation. However, it utilizes two diffusion models: one for learning CLIP embeddings and the other for using CLIP as a condition to generate images. It reveals that flow matching loss is more effective than MSE loss, enabling more diverse image sampling and yielding better image quality. UniLIP [172] incrementally incorporates reconstruction ability into CLIP through the self-distillation, then employs a learnable query along with the hidden state of the last layer of the MLLM as combined conditions. This framework is demonstrated to optimize the abundant information for visual editing. To exploit the hierarchical representations within the MLLM’s intermediate layers, TBAC-UniImage [173] applies learnable queries at multiple layers instead of the last layer. To sum up, these learnable query-based designs share a common strength: they provide adaptive, compact, and semantically enriched representations that support both efficient image understanding and high-quality generation. By focusing on task-driven token extraction, such models offer a flexible and extensible alternative to traditional visual tokenizers, especially in unified multimodal frameworks.
@@
Pseudo Hybrid Encoding. Representative works in this category include Janus [174], Janus-Pro [175], OmniMamba [176], Unifluid [177], and MindOmni [178]. As shown in Fig. 5 (b-4), these models adopt dual encoders—typically a semantic encoder (e.g., SigLIP) and a pixel encoder (e.g., VQGAN or VAE)—but use them in a task-specific manner. During training, the semantic encoder branch is enabled for vision-language understanding tasks, while the pixel encoder branch is activated for image generation tasks. Although the dual encoders are trained concurrently with combined understanding and generation datasets, the pixel encoder is not utilized during inference in understanding tasks and the semantic encoder is disabled for text-to-image generation. However, for image editing, Unifluid [177] uses the semantic encoder to encode the source image while MindOmni [178] utilizes both VAE and semantic encoder to encode the source image. The rationale behind this design choice is that mixed training with both types of data can enhance performance across understanding and generation tasks. Skywork UniPic [179] employs SigLIP2 as the encoder for understanding tasks and MAR [25] as the encoder for generative tasks. However, since only one encoder is active at any given time, these models do not fully harness the advantages of hybrid encoding. Specifically, they miss the opportunity to employ semantic grounding in generation tasks and fail to utilize high-fidelity visual details in comprehension tasks. Consequently, these models typically engage pixel decoders to reconstruct images from latent codes.
@@ 4.2 Text-to-Image Datasets
  • Synthesized Datasets: Specialized datasets for text-to-image synthesis are increasingly created using existing generative models. The text-to-image-2M dataset [261] provides  2 million enhanced text-image pairs for fine-tuning, curated using advanced T2I and captioning models. SFHQ-T2I [263] offers 122K diverse, high-resolution synthetic face images generated by multiple T2I models, ensuring variance and privacy. For entity control, the EliGen TrainSet [264] uses images from a baseline model (FLUX.1-dev) and MLLM-generated prompts for stylistic consistency and detailed annotation. Similarly, BLIP-3o 60k [169] provides 60,000 instruction tuning samples distilled from GPT-4o, covering various categories for diverse training. ShareGPT-4o-Image[266] contributes 45K text-to-image pairs, where prompts are generated through both a structured attribute-first approach and an image-first approach, with corresponding images synthesized by GPT-4o’s image generation capabilities to distill its advanced skills. To specifically address blind spots in real-world data, Echo-4o-Image [267] provides over 100K samples targeting surreal fantasy scenarios and complex, long-tail instructions to enhance model imagination and alignment.